# MNO: Multiscale Neural Operator for Computational Fluid Dynamics with 3D Point Cloud Data

### v1

@@ 4.4 Visualization of Attention Modules
Figure 2 visualizes the prediction errors for different attention configurations. Each row corresponds to one of the four benchmarks, while columns represent the module combinations: the first column shows predictions using only Global Attention; the second column shows Global+Local Attention; the third column repeats the second but with a different color scale for better contrast; and the fourth column shows the full combination of Global, Local, and Micro Attention. The following discussion takes Figure 2(b) as an illustrative example.
@@
Figure 2: The visualization of Global, Local and Micro Attention modules. The red circle serves as a reference for areas with obvious differences. Rows show the error maps for different benchmarks. For each row: (a) ShapeNet Car. The arrow represents the direction of the wind, and the color denotes the prediction error; (b) Ahmed body; (c) DrivAerNet++; (d) Parachute. For each column: (first column) prediction of only Global Attention module; (second column) prediction of Global and Local Attention modules together; (third column) identical values to the second column but with a different color scale; (fourth column) prediction of the full MNO with Global, Local, and Micro Attention modules together.
5 Conclusions
Appendix A LLMs Polishing
The manuscript was initially drafted in Chinese and polished using Large Language Models (LLMs) of DeepSeek-R1 and ChatGPT-4.0 to improve clarity, grammar, and academic style. The authors rigorously reviewed and edited all AI-generated content to ensure accuracy and consistency with the original scientific intent. The intellectual contributions remain entirely human.
@@ Appendix B Details of Benchmarks
This paper conducts a comprehensive evaluation of the model across four benchmarks. The details for each benchmark are provided below.
The ShapeNet Car (Umetani & Bickel, 2018) focuses on wind tunnel experiments for automobiles, a critical stage in automotive industrial design. This dataset contains 889 samples representing different car shapes, used to simulate driving conditions at a speed of 72 km/h. The car shapes are drawn from the ”Car” category of ShapeNet (Chang et al., 2015). The surrounding space is discretized into an unstructured grid with 32,186 points, and both the airflow velocity around the car and the pressure on the car surface are recorded. The number of points on the car surface is 3,682. Following the experimental setup in Transolver (Wu et al., 2024), we use 789 samples for training and the remaining 100 samples for testing. The input point cloud of each sample is preprocessed into a combination of point positions, signed distance functions, and normal vectors. A notable difference is that the original dataset contains 96 fixed noisy points on the car surface. After our preprocessing, the point cloud data consists of 29,498 air points and 3,586 car surface points.
@@
The Ahmed Body (Li et al., 2023b) is a wind tunnel dataset for bluff-body vehicles, used to predict the pressure on the vehicle surface. The vehicle shape is based on the benchmark model designed in (Ahmed et al., 1984). The inflow velocity ranges from 10 m/s to 70 m/s, corresponding to Reynolds numbers from $4.35\times 10^{5}$ to $6.82\times 10^{6}$. The dataset is generated by systematically varying the vehicle’s length, width, height, ground clearance, inclination angle, and rear rounding radius, resulting in a total of 551 samples, each containing approximately 100k surface points. Among these, 500 samples are used for training and 51 samples for testing, consistent with the setup in PCNO (Zeng et al., 2025).
@@
The Parachute Dynamics (Zeng et al., 2025) captures the inflation process of different parachutes under specific pressure loads. The pressure load increases linearly from 0 to 1000 Pa over the first 0.1 seconds and then remains constant at 1000 Pa. The learning objective is to map the initial parachute shape to the displacement fields at four specific time points during inflation: $t_{1}$ = 0.04, $t_{2}$ = 0.08, $t_{3}$ = 0.12, and $t_{4}$ = 0.16. These time points characterize the inflation process, where the parachute first rapidly expands under pressure, then over-expands, and finally rebounds. The experimental setup follows that of PCNO (Zeng et al., 2025), with 1000 samples for training and 200 samples for testing. Each sample contains approximately 15k points in the point cloud.
Appendix C The Evaluation Metrics
C.1 The Relative L2 Error
C.2 The Mean Absolute Error
Appendix D Full Implementation Details
@@
The implementation software of the model is mainly based on PyTorch 2.4.1, CUDA 12.1, and Python 3.9.0. The computing platform mainly includes Ubuntu 22.04.4 LTS and 4 NVIDIA H100 GPUs.
Table 4: The hyperparameters and training configuration of MNO for different benchmarks.
Benchmark N Block M D Head $k_{nbr}$ Batch Epoch Loss Optimizer Scheduler
ShapeNet Car 30k 4 256 128 8 16 4 500 RL2 AdamW OneCycleLR
DrivAerNet++ 300k 8
Ahmed body 100k 16
Parachute 15k 16
@@
Table 4 details the hyperparameter configurations used during training. Here, N denotes the scale of point clouds per sample, Block indicates the number of MNO blocks, M represents the number of tokens in the low rank space of the Global Attention module, D signifies the vector dimension per token, Head refers to the number of heads in the Multi-head Self-Attention (MSA) mechanism, and $k_{nbr}$ specifies the number of neighboring nodes in the Local Attention module. The training employed a relative L2 error loss function over 500 epochs. Optimization is performed using the AdamW optimizer coupled with a OneCycleLR learning rate scheduler, with the initial learning rate set to $1\times 10^{-3}$.
Appendix E The Ablation Experiments of the Depth of MNO
@@
Table 5 presents the experimental results. It is evident that the MNO model incorporating all three attention modules achieves the highest prediction accuracy in most cases. A significant improvement in MNO’s predictive performance is observed as the number of blocks increases from 1 to 4. However, performance gains become marginal when the block count exceeds 4, suggesting that the model likely enters an overfitting state at this stage.
Notably, the predictive performance of Local Attention and Micro Attention modules exhibits a positive correlation with the number of blocks, whereas Global Attention module performance shows an inverse correlation. This occurs because the low rank projection within Global Attention discards certain mid-to-high frequency information. Cascading multiple MNO blocks increases the number of low rank projections, thereby amplifying the loss of detailed information. In contrast, Local Attention incorporates $pos_{rel}$ to mitigate feature distortion caused by network deepening, and Micro Attention employs residual branches to supplement fine-grained details.
In summary, as the depth of the MNO model increases, the division of labor among different attention modules becomes more distinct. Global Attention increasingly focuses on analyzing global low-frequency features, while Local Attention and Micro Attention dedicate greater emphasis to capturing local mid-to-high-frequency details.
Appendix F The Display and Discussion of Prediction Results of MNO Model

### v2

@@ Appendix A Visualization of Attention Modules
Figure 3 visualizes the prediction errors for different attention configurations. Each row corresponds to one of the five benchmarks, while columns represent the module combinations: the first column shows predictions using only Global Attention; the second column shows Global+Local Attention; the third column repeats the second but with a different color scale for better contrast; and the fourth column shows the full combination of Global, Local, and Micro Attention. The following discussion takes Figure 3 (b) as an illustrative example.
Figure 3: The visualization of Global, Local and Micro Attention modules. The red circle serves as a reference for areas with obvious differences. Rows show the error maps for different benchmarks. For each row: (a) ShapeNet Car. The arrow represents the direction of the wind, and the color denotes the prediction error; (b) Ahmed body; (c) DrivAerNet++; (d) Magnetohydrodynamics. All fields are cross sections at the z-axis center; (e) Parachute. For each column: (first column) prediction of only Global Attention module; (second column) prediction of Global and Local Attention modules together; (third column) identical values to the second column but with a different color scale; (fourth column) prediction of the full MNO with Global, Local, and Micro Attention modules together.
@@ Appendix B Details of Benchmarks
This paper conducts a comprehensive evaluation of the model across five benchmarks. The details for each benchmark are provided below.
The Magnetohydrodynamics (MHD) (Ohana et al., 2024) is designed for astrophysical and plasma turbulence research. The ideal MHD equations are solved using a third-order accurate hybrid essentially non-oscillatory (ENO) scheme, with periodic boundary conditions, an isothermal equation of state, and random large-scale solenoidal forcing. Key parameters include the sonic Mach number ($M_{S}$ = 0.5, 0.7, 1.5, 2.0, 7.0) and Alfvén Mach number ($M_{A}$ = 0.7, 2.0). The dataset includes 3D data of density (scalar), velocity (vector), and magnetic field (vector), with 78 training and 10 test samples. Each sample has dimensions (time, x, y, z, fields) = (100, 64, 64, 64, 7), where fields: 1 (density) + 3 (velocity) + 3 (magnetic). The goal is to predict all fields at the next time step using observations from the previous four time steps. Each sample has approximately 260k points in point cloud representation.
The ShapeNet Car (Umetani and Bickel, 2018) focuses on wind tunnel experiments for automobiles, a critical stage in automotive industrial design. This dataset contains 889 samples representing different car shapes, used to simulate driving conditions at a speed of 72 km/h. The car shapes are drawn from the ”Car” category of ShapeNet (Chang et al., 2015). The surrounding space is discretized into an unstructured grid with 32,186 points, and both the airflow velocity around the car and the pressure on the car surface are recorded. The number of points on the car surface is 3,682. Following the experimental setup in Transolver (Wu et al., 2024), we use 789 samples for training and the remaining 100 samples for testing. The input point cloud of each sample is preprocessed into a combination of point positions, signed distance functions, and normal vectors. A notable difference is that the original dataset contains 96 fixed noisy points on the car surface. After our preprocessing, the point cloud data consists of 29,498 air points and 3,586 car surface points.
@@
The Ahmed Body (Li et al., 2023b) is a wind tunnel dataset for bluff-body vehicles, used to predict the pressure on the vehicle surface. The vehicle shape is based on the benchmark model designed in (Ahmed et al., 1984). The inflow velocity ranges from 10 m/s to 70 m/s, corresponding to Reynolds numbers from $4.35\times 10^{5}$ to $6.82\times 10^{6}$. The dataset is generated by systematically varying the vehicle’s length, width, height, ground clearance, inclination angle, and rear rounding radius, resulting in a total of 551 samples, each containing approximately 100k surface points. Among these, 500 samples are used for training and 51 samples for testing, consistent with the setup in PCNO (Zeng et al., 2025b). The variations in inflow velocity and Reynolds number are shown in Figure 4.
Figure 4: Sample distribution of varying inflow velocities and Reynolds numbers for the Ahmed body benchmark case.
@@
The Parachute Dynamics (Zeng et al., 2025b) captures the inflation process of different parachutes under specific pressure loads. The pressure load increases linearly from 0 to 1000 Pa over the first 0.1 seconds and then remains constant at 1000 Pa. The learning objective is to map the initial parachute shape to the displacement fields at four specific time points during inflation: $t_{1}$ = 0.04, $t_{2}$ = 0.08, $t_{3}$ = 0.12, and $t_{4}$ = 0.16. These time points characterize the inflation process, where the parachute first rapidly expands under pressure, then over-expands, and finally rebounds. The experimental setup follows that of PCNO (Zeng et al., 2025b), with 1000 samples for training and 200 samples for testing. Each sample contains approximately 15k points in the point cloud.
Appendix C Full Implementation Details
@@
The implementation software of the model is mainly based on PyTorch 2.4.1, CUDA 12.1, and Python 3.9.0. The computing platform mainly includes Ubuntu 22.04.4 LTS and 4 NVIDIA A800 GPUs.
Table 4: The key hyperparameters and training configurations of our MNO and baseline methods.
@@
Table 4 details the hyperparameter configurations and training settings of the MNO model and other baseline methods across all benchmarks. Training employed a relative L2 error loss function over 500 epochs, optimized using the AdamW optimizer and the OneCycleLR learning rate scheduler. In the MNO model, $N$ denotes the point number per sample, Block refers to the number of MNO blocks, $M$ indicates the number of tokens in the low rank space of the global attention module, $D$ represents the vector dimension of each token, $Head$ specifies the number of heads in the multi-head self-attention (MSA) mechanism, and $k_{nbr}$ defines the number of neighboring nodes in the local attention module. For MHD and DrivAerNet++, due to GPU memory limitations, $k_{nbr}$ is reduced to 8.
The parameter configurations for all baseline models are derived from the settings provided in the authors’ official papers and code repositories. In Transolver (Wu et al., 2024), n_layers is similar in meaning to Blocks in MNO, and slice_num represents the number of physical slices in the latent space. In LNO (Wang and Wang, 2024), n_block refers to the number of MSA modules in the latent space, and n_mode refers to the number of tokens in latent space. In PointNet++ (Qi et al., 2017b), point numbers refer to the quantity of spatially downsampled or upsampled points at different scale levels, and feature dims refer to the number of channels at each level. This resembles a U-Net-like architecture for point clouds. Other detailed parameter explanations will not be reiterated. Readers can refer to the baselines’ original papers and official code repositories for further details.
Appendix D Computational cost
In this section, we present a detailed analysis of the computational cost of MNO. For a fair comparison, the training and inference times of all models are measured on the same hardware platform (a single A800 GPU with 80 GB memory), with the batch size fixed to 1.
Table 5: Statistical results of computational costs for MNO model. FlopsFlops and $Params$ represent the model’s theoretical computational load and parameter count, respectively. $GPU_{t}$ and $GPU_{i}$ denote the training and inference GPU memory per sample, respectively. train time /epoch and inference time /sample stand for the training time per epoch and the inference time per sample, respectively. ”>80G” and ”OOM” denote out of memory.
Ahmed body (100k)
Modules Flops Params $GPU_{t}$ $GPU_{i}$ train time /epoch inference time /sample
Global (no low rank) 116.6G 1.16M >80G >80G OOM OOM
Global 96.5G 1.49M 7.1G 1.0G 483s 0.3489s
Global+Local 442.6G 1.95M 37.7G 8.4G 932s 1.2054s
Global+Local+Micro 462.5G 2.15M 38.6G 8.4G 939s 1.2218s
Parachute (15k)
Modules Flops Params $GPU_{t}$ $GPU_{i}$ train time /epoch inference time /sample
Global (no low rank) 16.759G 1.16M >80G 19.7G OOM 0.0946s
Global 13.9G 1.49M 1.8G 0.6G 120s 0.0432s
Global+Local 63.7G 1.95M 5.9G 1.67G 203s 0.1015s
Global+Local+Micro 66.5G 2.15M 6.1G 1.67G 205s 0.0979s
Magnetohydrodynamics (260k)
Modules Flops Params $GPU_{t}$ $GPU_{i}$ train time /epoch inference time /sample
Global (no low rank) OOM OOM >80G >80G OOM OOM
Global 252.297G 1.497M 17.2G 2.2G 185s 1.0702s
Global+Local 737.091G 1.963M 57.3G 11.8G 450s 4.2957s
Global+Local+Micro 788.765G 2.161M 59.5G 11.8G 453s 4.3174s
ShapeNet Car (30k)
Modules Flops Params $GPU_{t}$ $GPU_{i}$ train time /epoch inference time /sample
Global (no low rank) 37.2G 1.16M >80G >80G OOM OOM
Global 30.9G 1.49M 2.7G 0.7G 278s 0.1219s
Global+Local 141.4G 1.95M 12.4G 3.1G 402s 0.2917s
Global+Local+Micro 147.8G 2.15M 12.7G 3.1G 408s 0.2971s
DrivAerNet++ (300k)
Modules Flops Params $GPU_{t}$ $GPU_{i}$ train time /epoch inference time /sample
Global (no low rank) OOM OOM >80G >80G OOM OOM
Global 286.8G 1.49M 19.7G 2.4G 533s 1.2204s
Global+Local 841.6G 1.95M 65.4G 13.4G 1404s 5.3302s
Global+Local+Micro 900.8G 2.15M 67.9G 13.4G 1413s 5.3542s
In Table 5, ”Global (no low rank)” refers to the standard self-attention method, which computes global attention across all spatial points directly without using low rank projection for compression. Compared to this standard self-attention method, the low rank projection reduces the inference time by 54.3%. Furthermore, due to training GPU memory requirements exceeding 80GB, the standard self-attention method could not complete the training task on any benchmark used in this paper.
“Global+Local+Micro” denotes the full MNO model. For physical fields with 15k, 30k, 100k, 260k, and 300k points, MNO requires only 0.09s, 0.3s, 1.2s, 4.3s, and 5.3s for inference, respectively. This demonstrates second-level inference latency, meeting the requirements of large-scale CFD prediction. Moreover, MNO supports multi-sample parallelism, with further efficiency gains achieved by increasing the batch size or the number of GPUs.
Table 6: Comparison of computational costs with the latest baseline. ”Params” represent the model’s parameter count, respectively. “train/epoch“ and ”inference/sample“ are the training time per epoch and the inference time per sample, respectively. $n_{train}$ denotes the training samples of the benchmark. NN is the number of points per sample.
Benchmark NN $n_{train}$ Method Params train/epoch inference/sample
Parachute 15k 1000 AMG(Li et al., 2025) 2.15M 0.29h 1.0335s
Ours 2.15M 0.06h 0.0979s
ShapeNet Car 30k 789 AMG(Li et al., 2025) 2.15M 0.88h 3.9245s
Ours 2.15M 0.11h 0.2971s
Ahmed body 100k 500 AMG(Li et al., 2025) 2.15M 4.58h 32.828s
Ours 2.15M 0.26h 1.2218s
Magnetohydrodynamics 260k 78 AMG(Li et al., 2025) 2.15M 4.81h 221.54s
Ours 2.15M 0.13h 4.3174s
DrivAerNet++ 300k 200 AMG(Li et al., 2025) 2.15M 16.22h 291.15s
Ours 2.15M 0.39h 5.3542s
Table 6 compares the computational speed of MNO with the recent SOTA method AMG under comparable parameter budgets. Although both methods adopt graph structures and multi-scale strategies, MNO reduces actual inference time by about 90%. This is because AMG follows a conventional multi-level point-cloud reconstruction strategy, where the computational graph is rebuilt at each stage based on the current features, and adjacency relationships are computed using FPS. In contrast, all MNO blocks share a single graph, so adjacency is computed only once at the frontend, substantially reducing computation. Meanwhile, the shared graph built on original spatial coordinates explicitly enhances geometric representation in the latent space, leading to a 21.1% accuracy improvement compared to AMG (Figure 2(a)).
Appendix E The Ablation of the Depth of MNO
@@
Table 7 presents the experimental results. It is evident that the MNO model incorporating all three attention modules achieves the highest prediction accuracy in most cases. A significant improvement in MNO’s predictive performance is observed as the number of blocks increases from 1 to 4. However, performance gains become marginal when the block count exceeds 4, suggesting that the model likely enters a saturated state at this stage.
MNO achieves satisfactory performance with only four serially connected blocks. Because that the kNN graph–based Local Module provides strong inductive bias and tends to saturate with relatively shallow depth, unlike deep Transformers that rely heavily on stacking for expressivity. Figure 2 also confirms this: MNO can achieve better performance than Transolver with fewer block numbers.
Each module operates at distinct receptive field scales: the global module captures domain-level dependencies, while the local and micro modules provide fine-grained neighborhood and point-level information. Conversely, if the global and micro modules function effectively in isolation, it indicates architectural redundancy rather than an effective and specialized model.
Appendix F The Ablation of $M$ in low rank space
To investigate the size of the model’s demand for low rank space representation capacity, we conduct ablation experiments with parameter $M$. Table 8 shows ablation results for $M$. When $M>=256$, model performance saturates, indicating limited capacity requirements in the low rank space. For CFD tasks with limited computational resources, we recommend reducing $M$, as performance does not drop significantly.
Table 8: The low rank space $M$ ablation results on ShapeNet Car.
Methods $M$ $RL2_{p}$ $MAE_{p}$ $RL2_{v}$ $MAE_{v}$ FlopsFlops ParamsParams $GPU_{t}$ $GPU_{i}$
MNO 8 0.0607 1.4416 0.0190 0.0911 143.584G 2.02M 12.4G 3.05G
MNO 16 0.0606 1.4320 0.0197 0.0917 143.720G 2.03M 12.4G 3.05G
MNO 32 0.0598 1.3803 0.0180 0.0857 143.992G 2.04M 12.4G 3.05G
MNO 64 0.0599 1.4123 0.0187 0.0910 144.536G 2.05M 12.4G 3.05G
MNO 128 0.0607 1.4201 0.0188 0.0888 145.625G 2.08M 12.6G 3.05G
MNO 256 0.0597 1.3796 0.0178 0.0845 147.802G 2.15M 12.7G 3.05G
MNO 512 0.0601 1.3763 0.0177 0.0822 152.155G 2.28M 13.0G 3.12G
Increasing the value of $M$ does not lead to significant improvements. This indicates that the low rank space does not rely on a large dimensionality setting. We speculate that among objects of the same category (such as different vehicles in ShapeNet Car), the overall contours exhibit high structural similarity, primarily composed of low-frequency information, while the key factors affecting prediction accuracy depend more on mid- to high-frequency details. Therefore, a smaller low rank space is sufficient to capture the low-frequency features of the overall contours, whereas finer geometric structures are dominated by mid- to high-frequency components. This result suggests that in flow field prediction tasks, low-frequency information primarily serves an auxiliary global constraint role, while mid- to high-frequency information is more critical for recovering local details.
Appendix G The Ablation of $k_{nbr}$ of Local Graph attention
To investigate the model’s sensitivity to the number of neighbors, an ablation study on the $k_{nbr}$ is conducted. The experimental results are shown in Table 9. When $k>=16$, the performance of the MNO reaches saturation. The GPU memory is linearly related to the value of k.
Table 9: The ablation results of $k_{nbr}$ of Local graph attention on ShapeNet Car. $GPU_{t}$ and $GPU_{i}$ represent the training and inference GPU memory per sample, respectively.
Methods $k_{nbr}$ $RL2_{p}$ $MAE_{p}$ $RL2_{v}$ $MAE_{v}$ FlopsFlops ParametersParameters $GPU_{t}$ $GPU_{i}$
Local 2 0.1902 5.8130 0.0627 0.3134 32.987G 0.83M 2.9G 0.97G
Local 4 0.1154 3.8378 0.0488 0.2407 45.742G 0.83M 4.1G 1.35G
Local 8 0.0943 3.2675 0.0428 0.2230 71.252G 0.83M 6.4G 1.91G
Local 16 0.0832 2.9235 0.0399 0.2042 122.272G 0.83M 11.2G 3.05G
Local 32 0.0746 2.4595 0.0350 0.1780 224.311G 0.83M 21.1G 5.32G
MNO 2 0.0648 1.4644 0.0184 0.0860 58.517G 2.15M 4.7G 1.02G
MNO 4 0.0645 1.4378 0.0186 0.0861 71.272G 2.15M 5.6G 1.44G
MNO 8 0.0635 1.4210 0.0183 0.0855 96.782G 2.15M 7.8G 1.91G
MNO 16 0.0597 1.3796 0.0178 0.0845 147.802G 2.15M 12.7G 3.05G
MNO 32 0.0599 1.4160 0.0191 0.0935 249.841G 2.15M 22.3G 5.32G
When the Local Module independently handles the CFD task, the prediction error rapidly decreases as $k_{nbr}$ increases. However, when the MNO model performs the CFD prediction task, the reduction in prediction error with increasing $k_{nbr}$ is more gradual. The Local Module lacks a global perspective of the point cloud. Increasing $k_{nbr}$ effectively expands its receptive field, making the Local Module highly sensitive to changes in $k_{nbr}$. Nevertheless, a larger $k_{nbr}$ significantly increases GPU memory usage and computational complexity, so its value cannot be set excessively high. The MNO model possesses receptive fields at multiple scales. The global perspective provided by the Global Module alleviates the Local Module’s strong dependence on a large receptive field, allowing the Local Module to focus more on analyzing local features. Consequently, the prediction performance of MNO is less sensitive to variations in $k_{nbr}$.
Since GPU memory is exceptionally sensitive to $k_{nbr}$, we must minimize the number of neighbor nodes $k_{nbr}$ to avoid memory overflow when the dataset contains a large number of sampled points. Fortunately, the unique multi-scale structure design of MNO mitigates the heavy reliance of graph structures on computational resources. As shown in Table 9, when the RL2 error is also less than 0.08, compared to the Local Graph Attention method, MNO reduces computational Flops by 73.9% and GPU memory usage by 83.8%. Therefore, even when $k_{nbr}$ is reduced to prevent memory overflow, MNO can still achieve excellent prediction performance. For instance, in the DrivAerNet++ dataset used in this study, the point cloud scale of 300k forces us to reduce $k_{nbr}$ from 16 to 8, yet the prediction error remains lower than that of the best baseline.
Appendix H Zero-Shot Resolution Adaptation Study
In this experiment, we designed an adaptive resolution training strategy that enables the MNO model to support point cloud inputs with arbitrarily varying resolutions within a certain range in a zero-shot manner. The training method is illustrated in the following equation:
$\begin{split}mask&=\text{randmask}(sample\ rate),\\ X^{sample}_{in}&=X_{in}\odot mask,\\ X^{sample}_{out}&=\text{model}(X^{sample}_{in}),\\ Y^{sample}&=Y\odot mask,\\ L&=\text{Loss}(X^{sample}_{out},Y^{sample})\end{split}$ (7)
where sampleratesample\ rate represents the proportion of sampling points, and $sample\ rate\in(10\%,100\%)$, $mask\in\mathbb{R}^{N_{\max}\times 1}$, and $N_{\max}$ represents the maximum limit of input points. The function $\text{randmask}(\cdot)$ randomly sets $n$ positions in the mask to 1, where $n=\mathrm{round}(N\times sample\ rate)$, and sets the remaining positions to 0. $X_{in}\in\mathbb{R}^{N_{\max}\times f}$ represents the original point cloud input data, where $f$ is the number of input features. $X^{sample}_{out}$ corresponds to the predicted physical field data of the sampled points, $Y$ is the ground truth, and Loss is the loss function.
Table 10: Zero-Shot results across different resolutions on DrivAerNet++ benchmark. Sample rate refers to the proportion of spatial sampling points. $RL2_{p}$, $MAE_{p}$, $RMSE_{p}$ and $MSE_{p}$ denote the Relative L2 error, Mean Absolute Error, Root Mean Square Error, and Mean Square Error of the pressure field, respectively.
sample rate $n$ $RL2_{p}$ $MAE_{p}$ $RMSE_{p}$ $MSE_{p}$
10% 30k 0.1793 15.7110 27.9917 816.6555
20% 60k 0.1727 15.1267 26.9546 743.8958
30% 90k 0.1708 15.0152 26.6468 725.7269
40% 120k 0.1728 14.9423 27.0181 753.2448
50% 150k 0.1729 14.9048 26.9991 759.3806
60% 180k 0.1712 14.9021 26.7007 735.7252
70% 210k 0.1715 14.9023 26.7680 739.0064
80% 240k 0.1725 14.9164 26.9393 754.6640
90% 270k 0.1719 14.9203 26.8491 747.4118
100% 300k 0.1717 14.9379 26.8093 743.3867
During model training, the hyperparameter settings are consistent with Table 4. The results are shown in Table 10. Across 10 input resolutions (100% → 10%), MNO maintains stable performance with RL2 = 0.1755 ± 0.0038. This proves that even in severely sparse situations, it exhibits strong robustness and consistent physical state prediction for subsampling and density variations.
Appendix I The Geometric Generalization Study
In this section, we investigate the geometric generalization capability of the MNO. The model is trained exclusively on one category of geometric shapes and tested on another category to evaluate its performance.
The car shapes in the DrivAerNet++ benchmark are categorized into Sedan and Sport Utility Vehicle (SUV), as illustrated in Figure 5. Sedans typically feature a low-center-of-gravity, streamlined body design that emphasizes road-handling stability and high-speed cruising capability. In contrast, SUVs are characterized by higher ground clearance and a more boxy, taller body, providing a more spacious cargo area and enhanced off-road performance. Accordingly, we conduct ablative experiments to evaluate geometric generalization performance for the two vehicle types.
Figure 5: Display of different geometric shapes of Sedan and SUV.
The experimental results are summarized in Table 11. The predicted RL2 errors for all four cases are below 0.22, indicating that the MNO possesses a certain degree of geometric generalization ability. It demonstrates that the model can learn some universal physical laws from training cases of a single geometric type. When the model is trained solely on Sedan-type shapes, the predicted RL2 error is approximately 0.14 for Sedan cases, while the error increases to around 0.20 for SUV cases. The converse also holds true. This suggests that the geometric generalization capability of the MNO is limited. The geometric homogeneity and limited quantity of the training samples cause the model to overfit to some extent.
Table 11: Research results on zero-shot transfer of different geometric shapes.
Train Test $RL2_{p}$ $MAE_{p}$ $RMSE_{p}$ $MSE_{p}$ Train samples Test samples
Sedan Sedan 0.1468 14.0959 23.3102 173.2603 200 50
Sedan SUV 0.2020 19.6659 31.2958 290.4522 200 50
SUV SUV 0.1419 12.4015 22.1808 162.3201 200 50
SUV Sedan 0.2189 20.1559 34.6224 368.7184 200 50
Increasing the diversity of training samples can effectively alleviate this issue. For example, if the training set of the MNO includes both vehicle types, the RL2 error will decrease from 0.2 to 0.16. In the future, we will collect and construct a large-scale 3D CFD benchmark encompassing a wider variety of geometric types and a larger sample size to enhance the generalization capability of the MNO.
Appendix J The Display and Discussion of Prediction Results of MNO Model